OpenStudy (anonymous):

A person standing at the edge of a seaside cliff kicks a stone over the edge with a speed of 20 m/s. The cliff is 54 m above the water’s surface, as shown.With what speed does the stone strike the water? Answer in units of m/s

OpenStudy (anonymous):

Using h=0.5*g*t^2, where (h=-54m and g=-9.8m/s) you can solve for how long it remains in the air, t. Then use vf=vi+gt.

OpenStudy (anonymous):

use the formula [v^{2}=u^2+2ax] where v= final velocity, u=initial velocity, a=acceleration due to gravity, x=distance the answer is 38.18

OpenStudy (anonymous):

aniruddha, that would be correct if the person kicked the stone vertically down at 20m/s. But since that's not possible, we should probably assume that 20m/s is the speed in the horizontal direction and 0m/s is the speed in the vertical direction. Using the same equation you used, the answer would then be 32.53m/s downwards.